Problem 2-25 Calculating Short Sale Returns (LO4, CFA5) You just sold short 700 shares of wetscope, inc . a fledgling software firm, at S81 per share. You cover your short when the price hits S71.50 per share one year later If the company paid $0.35 per share in dividends over this period, what is your rate of return on the investment? Assume an initial margin of 50 percent (Do not round intermediate calculatione. Enter your answer as a percent rounded to 2 decimal places) Rate of return
Answer #1

ANSWER = Rate of return = 22.59%

Proceeds from short sale = 700($81) = $56,700

Initial margin deposit = $56,700(.50) = $28,350

Total assets = Total liabilities and equity = $56,700 + 28,350 = $85,050

Cost of covering short = 700($71.50) = $50,050

Account equity = $85,050 – 50,050 = $35,000

Cost of covering dividends = 700($0.35) = $245

Dollar profit = $35,000 – 28,350 – 245 = $6,405

Rate of return = $6,405 / $28,350 = 22.59%